OpenStudy (anonymous):

The formula for the amount with simple interest is A = P(1 + rt). Find r, if A = 224, P = 200, and t = 2.

OpenStudy (ajprincess):

jst plug in the given values. 224=200(1+r*2) 224=200(1+2r)

OpenStudy (anonymous):

ok then how do i get the r what should i do

OpenStudy (ajprincess):

224=200(1+2r) 224=200+400r Pls follow the steps 1.Subtract 200 both sides 2.Divide by 400 both sides

OpenStudy (anonymous):

r = 0.6 r = 0.12 r = 0.06 r = 0.012

OpenStudy (anonymous):

r=0.12?

OpenStudy (ajprincess):

nope pls check again

OpenStudy (anonymous):

r=0.6

OpenStudy (anonymous):

A = P (1 + rt) 224 = 200 (1 + 2r) 224 = 200 + 400r 224 - 200 = 400r 24 = 400r 24/400 = r .06 = r

OpenStudy (ajprincess):